Three particles of masses 1.0kg,2.0kg and 3.0kg are placed at the corners A,B and C respectively of an equilateral triangle ABC of edge 1 m.Locate the centre of mass of the system.

Grade:11

4 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
14 years ago

Dear Nitin Vomani,

Ans:- Let the trangle ABC  is placed in such a way in the co ord plane so that the mid point of the side BC is the origin

point A=1kg

point B=2kg

point C=3kg

hence the X co ord is=[(2×(-1/2)+(3×1/2)+(1×0)]/(2+3+1)=1/12

The Y co ord is=[(2×0)+(3×0)+(1×√3/2)]/(1+2+3)=√3/12

Hence the COM is At=( 1/12  , √3/12)

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Nitin Vomani !!!

 


Regards,

Askiitians Experts
Soumyajit Das IIT Kharagpur

amartya
9 years ago
Sir the answer given by u does not match the answer given in the book the Xcm is right but the Ycm is given as root 3 /4 
prateek sonker
7 years ago
To find the center of mass:Fix one vertex of triangle as origin and give coordinates to other 2 vertices.apply the center of mass formulaIn the figure shown above A(0, 0) , B (1,0) and C(1/2, sqrt(3)/2)Xcm = M1X1+M2X2+M3X3/(M1+M2+M3) = 1(0)+2(1)+3(1/2)/(1+2+3) = 7/12Ycm = M1Y1+M2Y2+M3Y3/(M1+M2+M3) = 1(0)+2(0)+3(sqrt(3)/2)/(1+2+3) = sqrt(3)/4
Aman
6 years ago
A(0,0),B(1,0),C(1/2,√3/2)
Now as com at x=m1x1+m2x2+m3x3/m1+m2+m3 
=1(0)+2(1)+3(1/2)/1+2+3=(2+3/2)6=4+3/12=7/12 
com at y=m1y1+m2y2+m3y3/m1+m2+m3
              =1(0)+2(0)+3(√3/2)/1+2+3=(3√3/2)/6=3√3/2*6=√3/4
COM=√x*x+y*y=√7/12*7/12+√3/4*√3/4
           =√49/144+3/16
           =√49+27/144
            =√76/144
            =2√19/12 
          =   √19/6 m from A